OpenStudy (wasiqss):

determine the equation of parabola with an axis parralel to y axis, vertex on x axis,,n which passes through point(2,3) and (-1,12)

OpenStudy (anonymous):

This is easy man, what have you tried?

OpenStudy (wasiqss):

yes bt i dun have answers im getting focus (11/5) check if u get it

OpenStudy (anonymous):

i got it it's easy

OpenStudy (wasiqss):

yehh i want to recheck my answer

OpenStudy (anonymous):

wait i'm tapping...

OpenStudy (anonymous):

presume y=ax^2+bx^2+c so b^2-4ac<=0

OpenStudy (wasiqss):

lol temme answer,

OpenStudy (anonymous):

Sorry i went out

OpenStudy (anonymous):

i'm back !! a=-3 or -1/2, b=0 or -5/2 c=15 or10

OpenStudy (wasiqss):

how did u get it

OpenStudy (anonymous):

so y=-3x^2+15 or -1/2x^2-5/2+10

OpenStudy (wasiqss):

i mean how u took out focus, co ordinates

OpenStudy (anonymous):

haha ~ presume y=k(x-b)^2+d(dismiss d since the vertexis on the x axis){another way of the equation of parabola },and use the equation of parabola above at the same time,you try it

OpenStudy (anonymous):

okay with it?

OpenStudy (anonymous):

grait job explaining it all frank

OpenStudy (anonymous):

Thanks~